# PO-Flow: Flow-based Generative Models for Sampling Potential Outcomes and Counterfactuals

### v1

@@ 4.1 PO-Flow for Potential Outcomes Prediction
The potential outcomes prediction in PO-Flow follows a simple reverse process, as defined in Equation (11). Starting from a noise sample $y_{1}\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$. The procedure is summarized in Algorithm 1.
In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the probability densities of the potential outcomes, as detailed in Section 4.3.
@@ 4.2 PO-Flow for Counterfactual Outcomes Prediction
  Input: $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
  1. Forward process (encoding):
[…]
@@ 1. Forward process (encoding):
  3. Uncertainty-aware perturbations:
  4.     $\delta\sim N(0,\Sigma(x))$
  5.     $z^{\prime}_{y}=z_{y}+\delta$
  6. Reverse process (decoding):
[…]
         with $y(1)=z^{\prime}_{y}$.
@@ 6. Reverse process (decoding):
Under Assumption 3.2, the counterfactual outcome prediction in PO-Flow consists of both a forward and a reverse process. Given a factual observation $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 2–3 in Algorithm 2) encodes the information $(y^{(a)},x,a)$ into a latent representation centered at $z_{y}$. Lines 4–6 introduce uncertainty-aware perturbations to $z_{y}$, reflecting covariate-dependent uncertainty. Finally, the reverse process (Lines 7–8) decodes the latent $z_{y}^{\prime}$ back into the counterfactual outcome prediction $\hat{y}_{\text{cf}}^{(1-a)}$.
@@
Because of the smooth and information-preserving nature of the PO-Flow path trained via flow matching, the encoded $z_{y}$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$. To our knowledge, this is the first work to explore counterfactual outcome prediction using flow or diffusion-based models.
As illustrated in Figure 1, diffusion models rely on noisy and stochastic forward processes, which degrade the information in $z_{y}$ and lead to reversed samples that resemble interventional potential outcomes rather than counterfactuals. In Section 5.4, we present counterfactuals using an adapted implementation of DiffPO [43] (originally not designed for counterfactual), compared with PO-Flow.
@@ 4.3 PO-Flow for Learning Potential Outcomes Densities
Another distinctive advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs). We define $\hat{p}_{Y}(\cdot\mid x,a,y_{1})$ as the PO density learned by PO-Flow, conditioned on $(x,a)$ and a base sample $y_{1}$. The explicit form of the learned PO density is presented in Proposition 4.2, with the proof provided in Appendix A.
@@
Given a noise sample $y_{1}\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log\hat{p}_{Y}(\cdot|x,a,y_{1})=\log q(y_{1})+\int_{0}^{1}\nabla\cdot v_{\theta}(y(t),t,x,a)ds.$ (12)
Empirically, the divergence of the velocity field can be computed under the Hutchinson trace estimator (e.g., [31, 64]). More details are provided in Appendix B.5.
where $p(\epsilon)$ is a distribution in $\mathbb{R}^{d_{Y}}$ satisfying $\mathbb{E}[\epsilon]=0$ and Cov$(\epsilon)=I$ (e.g., a standard Gaussian).
To our knowledge, PO-Flow is the first generative approach in causal inference to explicitly learn the probability density of predicted potential outcomes. Unlike prior works [44, 45] that generate PO samples and estimate empirical distributions, PO-Flow provides direct access to likelihoods, enabling deeper insights into uncertainty.
This density estimation allows PO-Flow to assess prediction confidence via log-likelihood, identify unlikely outcomes, and quantify uncertainty. For counterfactual prediction, one can perturb the latent code $z_{y}^{\prime}$ multiple times and select the most confident $\hat{y}_{\text{cf}}^{(1-a)}$ based on log-likelihood, effectively reducing variance while improving robustness.
@@ 5 Experiments
Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner and T-learner [37]), representation-based conditional regressors (Counterfactual Regression (CFR) [55]), generative approaches (Causal Effect Variational Autoencoder (CEVAE) [42]; GAN for Individualized Treatment Effect (GANITE) [67]), and most recent generative approaches based on diffusion (DiffPO [43]) and discrete normalizing flows (INFs [45]).
Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1. Accordingly, we evaluate PO-Flow on standard synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
The code and datasets are available at https://github.com/StatFusion/PO-Flow.
@@ 5.1 Prediction of Potential Outcomes
[…]
@@
S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $1.01_{\pm.17}$ $1.16_{\pm.19}$
@@
Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}$ and the true outcomes $y_{i}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$. $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
@@ 5.2 Estimation of CATEs
Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y_{1}-y_{0}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. It captures individual-level heterogeneity in causal effects.
@@
Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets. Results are averaged over 10-fold cross-validation.
@@
S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $0.93_{\pm.08}$ $0.06_{\pm.02}$ $0.07_{\pm.04}$
@@ 5.3 Learning the Full Potential Outcome Distributions
Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the probability density of any generated outcome, providing deeper insights into predictive uncertainty.
@@ 5.3.1 Learning potential outcome densities
PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.2. To our knowledge, it is the first generative method in causal potential outcomes to provide this capability.
Evaluation Metrics: To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $\hat{p}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$. Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim\hat{p}}\left[\log\frac{\hat{p}(\hat{y}|x,a)}{\mathcal{N}(\mu(\hat{y}),1)}\right]$, where $\hat{p}$ is computed via (12). Table 3 reports the average results.
@@
This property allows PO-Flow to assess prediction confidence via log-density and to identify unlikely outcomes. For counterfactual estimation, one can sample multiple candidates from a perturbed latent region and select the most likely one based on the model’s estimated density.
@@ 5.3.2 Empirical distribution estimations
Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines. These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(\hat{p}(y|x,a),p(y|x,a))=\inf_{\gamma\in\Gamma(\hat{p},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$. The results are shown in Table 4.
@@
DiffPO $0.47_{\pm.17}$ $0.51_{\pm.18}$ $\mathbf{0.04_{\pm.02}}$ $\mathbf{0.13_{\pm.05}}$ $0.80_{\pm.15}$ $0.88_{\pm.17}$ $0.69_{\pm.14}$ $0.75_{\pm.14}$
INFs $0.59_{\pm.13}$ $0.67_{\pm.14}$ $0.08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
@@
Figure 2: Comparison between empirically estimated and the ground-truth POs distributions. Left: $P(y|x,a=0)$ on IHDP. Right: $P(y|x,a=1)$ on IHDP.
Understanding the distribution of potential outcomes (POs) is essential for reasoning about uncertainty in treatment decisions. As shown in Figure 2, PO distributions can shift differently across treatments for the same individual. PO-Flow provides not only point estimates but also distributional insight, supporting uncertainty-aware analysis. Section 5.3.1 further demonstrates its ability to estimate densities accurately. This enables practitioners to assess whether an outcome lies within a desired range and evaluate its likelihood (Proposition 4.2) under a given treatment, offering a clearer sense of confidence in the effect.

### v2

@@ 4.1 PO-Flow for Potential Outcomes Prediction
To sample potential outcomes $\hat{y}^{(a)}$ for an individual $(x,a)$, the procedure follows a simple reverse process, as defined in Equation (11). Starting from a noise sample $z\sim q(\cdot)$, we integrate the velocity field $v_{\theta}$ backward in time, conditioned on covariates $x$ and treatment $a$, to obtain a point estimate of the potential outcome $\hat{y}^{(a)}$. The steps are summarized in Algorithm 1.
In addition to predicting point estimates of potential outcomes, PO-Flow is capable of learning the densities of the potential outcomes, as detailed in Section 4.3.
@@ 4.2 PO-Flow for Counterfactual Outcomes Prediction
The counterfactual outcome prediction consists of both a forward and a reverse process. Given a factual $(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$, the forward process (Lines 1–2 in Alg. 2) encodes $(y^{(a)},x,a)$ into a latent representation $z$. Subsequently, the reverse process (Lines 3–5) decodes $z$ under the counterfactual treatment 1a1-a, yielding the counterfactual prediction $\hat{y}_{\text{cf}}^{(1-a)}$. The process of counterfactual prediction is also illustrated in Figure 1.
@@
Input: Factual sample $\small(y^{(a)},x,a)\sim p_{Y,X,A}(\cdot)$
 1. Forward process {abduction}:
3 2.  $z=y^{(a)}+\int_{0}^{1}v_{\theta}(y(t),t;x,a)\,dt,$
@@ 1. Forward process {abduction}:
 3. Reverse process {action and prediction}:
 4.  Set intervened action to 1a1-a;
4 5.  $\hat{y}_{0}=z-\int_{0}^{1}v_{\theta}(y(t),t;x,1-a)\,dt,$
   with $y(1)=z$.
@@ 4.  Set intervened action to 1a1-a;
Because of the smooth and information-preserving nature of the continuous normalizing flow path, the encoded $z$ retains rich information from the factual outcome $y^{(a)}$, enabling accurate decoding of the counterfactual $\hat{y}_{\text{cf}}^{(1-a)}$ under counterfactual treatment 1a1-a. To our knowledge, this work is among the early attempts at counterfactual prediction in the Potential Outcomes (POs) literature.
In the following, we present a supporting result on the counterfactual recovery properties of PO-Flow under monotone structural causal models (SCMs). While PO-Flow does not explicitly aim to model the true SCM during training or inference, we show that if the data are generated from monotone SCMs, then PO-Flow provides recovery guarantees under certain conditions.
We denote the SCM for an individual as $Y^{(A)}=f_{X}(A,U)$, where $U$ denotes the exogenous noise. The encoding function is given by $Z:=\Phi_{\theta}(Y^{(A)};\,X,A)$, which corresponds to Equation (10). We begin by introducing the following assumptions:
Assumption 4.1.
(A1) $U\perp\!\!\!\perp A$.
(A2) The structural causal model $f_{X}(A,U)$ is monotone in $U$ for each action $A\in\{0,1\}$.
(A3) For every $x$, the encoded latent $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ induced by the flow is statistically independent of $A$, i.e., $Z\perp\!\!\!\perp A\ |\ X=x$.
Remark 4.2.
In the POs framework, covariates $x$ are multivariate while the potential outcome is univariate, $y\in\mathbb{R}$. Therefore, (A2) is invoked for the univariate case. It is automatically satisfied under an additive structural causal model, $Y^{(A)}=f_{X}(A,U):=f_{X}^{*}(A)+U$. Under certain identifiability conditions, it also holds for non-linear models [62, 55]. The following theoretical results hold only under monotone SCM settings.
(A3) assumes the encoded representation $Z=\Phi_{\theta}(Y^{(A)};\,X,A)$ in PO-Flow is statistically independent of treatment $A$, though not necessarily functionally invariant to $A$. In the infinite-data limit with exact training, the continuous normalizing flow (Equation (8)) maps each treatment-conditioned outcome $Y$ to the same base distribution $q(Z)=\mathcal{N}(0,1)$, so $p_{\theta}(Z\mid X,A)=q(Z)$ and (A3) holds exactly. In finite samples, deviations may arise; we assess this empirically in Appendix D.1.
Under the Assumption 4.1, we present a result for PO-Flow when the true SCM is monotone:
Proposition 4.3 (Encoded $Z$ as a function of the exogenous noise $U$).
Let Assumption 4.1 hold. Assume that the exogenous noise $U\sim\mathrm{Unif}[0,1]$. Then there exists a continuously differentiable bijection function $\psi_{X}:\;\mathcal{U}\;\longrightarrow\;\mathcal{Z}$ that is independent of the treatment assignment $A$, i.e.,
$\small Z\;=\;\Phi_{\theta}(Y^{(A)};X,A)\;=\;\Phi_{\theta}(f_{X}(A,U);X,A)\;=\;\psi_{X}\bigl(U\bigr)$ (12)
Remark 4.4.
We can transform the exogenous noise $U\sim\mathrm{Unif}[0,1]$ assumed in Proposition 12 into other distributions, such as Gaussian. If $Z\sim\mathcal{N}(0,1)$ with CDF $F(Z)$, then $U=F(Z)\sim\mathrm{Unif}[0,1]$, and any structural assignment $f(\cdot,U)$ can be written equivalently as $\tilde{f}(\cdot,Z)=f(\cdot,F(Z))$. Therefore, the conclusions of Proposition 12 also hold with other noise distributions.
Following Proposition 12, we have the following result on counterfactual recovery under monotone SCMs:
Corollary 4.5 (Counterfactual recovery under monotone SCMs).
Assume the conditions in Assumption 4.1 hold. Consider a factual sample $\bigl(Y^{(A)},X,A\bigr)$ given by the structural equation $Y^{(A)}=f_{X}\bigl(A,U\bigr)$, and let its encoded representation be $Z=\Phi_{\theta}(Y^{(A)};X,A)$. Apply the intervention $\mathrm{do}\bigl(A=\gamma\bigr)$. Assume that the true counterfactual is given by $Y_{\mathrm{CF}}^{(\gamma)}=f_{X}(\gamma,U)$. The decoder then recovers the counterfactual almost surely:
$\hat{Y}_{\mathrm{CF}}^{\gamma}:=\Phi_{\theta}^{-1}\bigl(Z;\ X,\gamma\bigr)=Y_{\mathrm{CF}}^{\gamma}.$ (13)
Remark 4.6.
One may notice that Assumptions (A1) and (A2) mirror those in Bijective Generation Mechanisms (BGM) [43]. BGM establishes class-level identifiability: if a learned mechanism lies in the BGM class and matches the observational distribution ($\hat{f}_{X}(A,U)\stackrel{{\scriptstyle d}}{{=}}f_{X}(A,U)$), then counterfactuals are identifiable. By contrast, our Prop. 4.3 shows that PO-Flow’s encoder is bijective in $U$, placing our model within the BGM class. We then use an alternative constructive route (building on the intermediate step in Prop. 4.3) to obtain Cor. 4.5 under (A1)–(A3). Alternatively, one could invoke the BGM result after Prop. 4.3 to reach Cor. 4.5; see Appendix C.
@@ 4.3 PO-Flow for Learning Potential Outcomes Densities
Another advantage of PO-Flow is its ability to learn the density of the Potential Outcomes (POs). We define $p_{\theta,Y|X,A}(\cdot\mid x,a,z)$ as the PO density computed by PO-Flow, conditioned on $(x,a)$ and a base sample $z$. The explicit form of this density, mapping from latent to data space, is presented in Proposition 4.7, with the proof provided in Appendix A.
@@
Given a noise sample $z\sim q(\cdot)$ at $t=1$, the log-density of the resulting potential outcome at $t=0$, obtained via PO-Flow, is given by:
$\log p_{\theta,Y\mid X,A}\bigl(y(0)\mid x,a,z\bigr)=\log q(z)\;+\;\int_{0}^{1}\nabla_{y}\!\cdot\,v_{\theta}\bigl(y(t),t;\,x,a\bigr)\,dt.$ (14)
Empirically, the divergence of the velocity field can be computed much more efficiently under the Hutchinson trace estimator (e.g., [27, 58]). More details are provided in Appendix B.4.
Unlike prior works [40, 41] that generate PO samples and approximate empirical distributions, PO-Flow provides direct likelihood evaluation. This allows principled uncertainty quantification, supports ranking outcomes by plausibility (Table 4), enables information-theoretic analyses such as KL divergence (Table 3), and facilitates tasks like anomaly detection via quantile-based thresholds.
@@ 5 Experiments
Baselines: Our baselines cover several categories: model-agnostic learners that learn the conditional distribution means (S-learner [33], T-learner [33], and DR-Learner [32]), representation-based conditional regressors (Counterfactual Regression (CFR) [50]), deep generative methods (Causal Effect Variational Autoencoder (CEVAE) [38]; GAN for Individualized Treatment Effect (GANITE) [61]), and most recent generative approaches based on diffusion (DiffPO [39]) and discrete normalizing flows (INFs [41]).
Datasets: Counterfactual outcomes and unobserved potential outcomes (POs) are inaccessible in real-world data but can be simulated in semi-synthetic settings, as discussed in detail in Appendix B.1. Accordingly, we evaluate PO-Flow on synthetic benchmarks, including ACIC 2016, ACIC 2018, IHDP, and IBM.
The code and datasets are included in the supplementary materials for anonymous submission.
@@ 5.1 Prediction of Potential Outcomes
PO-Flow $\mathbf{0.53_{\pm.10}}$ $0.60_{\pm.12}$ $\mathbf{0.98_{\pm.11}}$ $\mathbf{1.19_{\pm.13}}$ $\mathbf{0.90_{\pm.09}}$ $\mathbf{0.97_{\pm.11}}$
@@
S-learner $\mathbf{0.54_{\pm.09}}$ $\mathbf{0.57_{\pm.11}}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
@@
Evaluation Metrics: We evaluate the point estimates of potential outcomes (POs) using the root mean squared error (RMSE) between the estimated outcomes $\hat{y}_{i}^{(a)}$ and the true outcomes $y_{i}^{(a)}$, computed as $\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}^{(a)}-\hat{y}_{i}^{(a)})^{2}}$, where $\text{RMSE}_{\text{in}}$ and $\text{RMSE}_{\text{out}}$ refer to the RMSE evaluated on the training and test sets, respectively.
@@
Besides, we also report in Table 4 the RMSE of the most likely POs selected by PO-Flow using the highest log-density (as computed in Proposition 4.7).
@@ 5.2 Estimation of CATEs
Conditional Average Treatment Effect (CATE): The CATE is defined as $\tau(x)=\mathbb{E}[y^{(a=1)}-y^{(a=0)}\mid X=x]$, representing the expected treatment effect conditioned on covariates $x$. It captures individual-level heterogeneity in causal effects.
@@
Table 2: Root Precision in Estimation of Heterogeneous Effect ($\sqrt{\epsilon_{\text{PEHE}}}$) across different methods on the ACIC 2016, ACIC 2018, IHDP, and IBM datasets.
@@
PO-Flow $\mathbf{0.76_{\pm.08}}$ $\mathbf{0.82_{\pm.09}}$ $\mathbf{0.05_{\pm.01}}$ $\mathbf{0.07_{\pm.02}}$ $\mathbf{0.41_{\pm.08}}$ $\mathbf{0.45_{\pm.10}}$ $\mathbf{0.04_{\pm.01}}$ $\mathbf{0.04_{\pm.01}}$
@@
S-learner $3.81_{\pm.76}$ $3.87_{\pm.98}$ $\mathbf{0.06_{\pm.02}}$ $\mathbf{0.07_{\pm.03}}$ $0.92_{\pm.07}$ $93_{\pm.08} 0.0$$6_{\pm.02} 0.0$$07_{\pm.04}$
@@ 5.3 Learning the Full Potential Outcome Distributions
Beyond point estimation, PO-Flow offers two distinct advantages for modeling potential outcomes: (1) it enables individualized sampling from the conditional distribution $P(Y\mid X,A)$, generating a rich set of plausible outcomes for each individual; and (2) it supports learning the density of any generated outcome, providing deeper insights into predictive uncertainty.
@@ 5.3.1 Learning potential outcome densities
PO-Flow enables direct computation of potential outcome densities, as shown in Proposition 4.7. To evaluate the learned densities, we compute the KL divergence between the PO-Flow estimate $p_{\theta,Y|X,A}(\cdot|x,a)$ and the ground-truth density $p_{\text{true}}(\cdot|x,a)$. Given that the semi-synthetic datasets assume $y^{(a)}\sim\mathcal{N}(\mu(a),1)$, the individual-level KL is $\text{KL}(\hat{Y}_{X}^{(a)}\|Y_{X}^{(a)})=\mathbb{E}_{\hat{y}\sim p_{\theta,Y|X,A}}\left[\log\frac{p_{\theta,Y|X,A}(\hat{y}|x,a)}{\mathcal{N}(\hat{y};\mu(a),1)}\right]$, where $p_{\theta,Y\mid X,A}$ is computed via (14).
@@
This property enables PO-Flow to identify most likely potential outcomes based on their log-densities. Table 4 reports the RMSE of POs selected by PO-Flow with the highest log-density, illustrating one application of this property.
Table 4: Root mean squared error (RMSE) of estimated potential outcomes (POs) across different methods on the ACIC 2018, IHDP, and IBM datasets. For PO-Flow, results are based on the most likely PO selected from 100 simulated outcomes using log-density.
ACIC 2018 IHDP IBM
PO-Flow $\mathbf{0.42_{\pm.07}}$ $\mathbf{0.54_{\pm.10}}$ $\mathbf{0.96_{\pm.09}}$ $\mathbf{1.05_{\pm.10}}$ $\mathbf{0.84_{\pm.06}}$ $\mathbf{0.90_{\pm.08}}$
INFs $0.69_{\pm.16}$ $0.78_{\pm.17}$ $1.02_{\pm.10}$ $1.20_{\pm.12}$ $1.52_{\pm.24}$ $1.57_{\pm.28}$
S-learner $0.54_{\pm.09}$ $0.57_{\pm.11}$ $1.31_{\pm.18}$ $1.44_{\pm.20}$ $01_{\pm.17} 1.1$$16_{\pm.19}$
CEVAE $0.83_{\pm.17}$ $0.85_{\pm.17}$ $1.18_{\pm.15}$ $1.36_{\pm.18}$ $0.96_{\pm.12}$ $0.98_{\pm.12}$
CFR $0.94_{\pm.10}$ $0.98_{\pm.12}$ $10_{\pm{.20}} 1.1$$18_{\pm{.20}}$ $2.09_{\pm.45}$ $2.17_{\pm.48}$
Compared with the averaged RMSE over all POs generated by PO-Flow in Table 1, selecting the most likely POs based on log-density yields smaller RMSEs, providing more confident and closer estimates.
@@ 5.3.2 Empirical distribution estimations
Evaluation Metrics: To enable comparison of estimated potential outcome distributions across methods, we include empirical distribution estimates for all baselines. These are evaluated using the Wasserstein-1 distance, defined as: $W_{1}(p_{\theta,Y|X,A}(y|x,a),p_{Y|X,A}(y|x,a))=\inf_{\gamma\in\Gamma(p_{\theta},p)}\mathbb{E}_{(\hat{y},y\sim\gamma)}[\|\hat{y}-y\|_{1}]$. The results are shown in Table 5.
@@
PO-Flow $\mathbf{0.36_{\pm.11}}$ $\mathbf{0.42_{\pm.12}}$ $\mathbf{0.05_{\pm.03}}$ $\mathbf{0.12_{\pm.04}}$ $\mathbf{0.30_{\pm.05}}$ $\mathbf{0.41_{\pm.06}}$ $\mathbf{0.33_{\pm.07}}$ $\mathbf{0.44_{\pm.08}}$
@@
INFs $0.59_{\pm.13}$ $67_{\pm.14} 0.0$$08_{\pm.03}$ $0.16_{\pm.05}$ $0.32_{\pm.08}$ $\mathbf{0.40_{\pm.09}}$ $0.36_{\pm.08}$ $\mathbf{0.45_{\pm.09}}$
@@
Understanding the distribution of potential outcomes (POs) is essential for uncertainty-aware treatment decisions, as outcomes for the same individual may vary substantially across treatments. PO-Flow provides not only point estimates but also full distributional insight, enabling practitioners to assess whether outcomes fall within desired ranges and to evaluate their likelihood (Proposition 4.7).